OpenStudy (anonymous):

would someone be so kind to show me the steps for solveing 1/sqrt2.sqrt3/2 - 1/2.1\sqrt2

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2} }\times \frac{ \sqrt{3} }{ 2 }-\frac{ 1 }{ 2 }\times \frac{ 1 }{ \sqrt{2} }\]

OpenStudy (anonymous):

just so its a little cleaner this is correct right

OpenStudy (anonymous):

thanks i just had to jolt my memory sqrt3-1 /2sqrt2

OpenStudy (anonymous):

most teachers dont want a square root as the denominator